Mean equal to 15 and standard deviation equal to 5. Find the probabilities associated with getting:
a. A 15 or above
b. Between a 15 and an 18
c. A 13 and above

Answer by stanbon(48568)

Mean equal to 15 and standard deviation equal to 5. Find the probabilities associated with getting:
a. A 15 or above
Since 15 is the mean,P(15 or above) = 1/2 or 0.50
---------------
b. Between a 15 and an 18
15 has a z-score of 0; 18 has a z-score of (18-15)/5 = 3/5=0.6
Using a z chart P(0 ----------------------
c. A 13 and above
P(z>0.6) = 0.50001-0.2257 = 0.2743
---------------------
Cheers,
Stan H.